What will be the final temperature if 50g of water at exactly 0°C is added to 250g of water at 90°C?


1
Expert's answer
2022-04-01T16:32:57-0400

Given:

"t_1=0^\circ\rm C"

"t_2=90^\circ\rm C"

"m_1=50\:\rm g"

"m_2=250\:\rm g"

The heat balance gives


"Q_1=Q_2\\
cm_1(t_3-t_1)=cm_2(t_2-t_3)"

Hence, the final temperature

"t_3=\frac{m_1t_1+m_2t_2}{m_1+m_2}"

"t_3=\frac{50*0+250*90}{50+250}=75^\circ\rm C"